If a and b are positive integers and a^2-b^2 = 7 what is the value of a ?
A 3 b 4 c 5 d 6 e 7

Answer by stanbon(57967)

If a and b are positive integers and a^2-b^2 = 7 what is the value of a ?
------------------
b cannot be 1 or 2 because "a" would not be an integer.
----------------------------
The smallest value "b" can have is 3.
Then a^2-9 = 7
a^2 = 16
a = 4
-----------------
Cheers,
Stan H.
=================
A 3 b 4 c 5 d 6 e 7